OpenStudy (ivyjane):

Can you divide 4x by 2? or does it have to be a 2x

OpenStudy (jiteshmeghwal9):

yes u may diide 4x by 2 & u gt 2x in rsult

OpenStudy (phi):

what do you mean ? you can always write \[ \frac{4x}{2} \] which means "4x dvided by 2" and you can simplify if you want to 2x (usually people like simple, but it does not have to be simplified)

OpenStudy (phi):

you can also think of 4x/2 as \[ 4 \cdot x \cdot \frac{1}{2} \] (dividing by 2 is the same as multiplying by 1/2) when you multiply, you can change the order, so you can write it as \[ \frac{1}{2} \cdot 4 \cdot x\] when multiplying, you can do the first two i.e. do 1/2 * 4 which is 2 and get 2⋅x people usually leave out the multiply sign (the "dot" here) and just write 2x

OpenStudy (mathmale):

What you should do here depends upon the context. Somequestions may ask you to divide 2x by 2; others may ask you to divide 2x by 2x. What are the respective quotients (answers)?